OpenStudy (anonymous):

Find the area of the following circle. (Round answer to the nearest hundredth). r = 4.5 cm Area = _____ cm² Circumference = ____ cm

OpenStudy (anonymous):

@JakeV8

OpenStudy (anonymous):

what is the formula for area of circle?

OpenStudy (anonymous):

area=pi*r^2=63.58 cm^2 cf=2*pi*r=28.26 cm

OpenStudy (anonymous):

hey, there it is!!

OpenStudy (anonymous):

PIxR^2

OpenStudy (anonymous):

right... and you know that r = 4.5 cm So put r = 4.5 cm into that formula. But the problem says to round the answer to hundredths... 2 decimal positions.

OpenStudy (anonymous):

a=pix20.25

OpenStudy (anonymous):

i got 63.61

OpenStudy (anonymous):

if you have a calculator with pi, you can multiply that out using the "pi" button. Otherwise, you can estimate pi as 3.14159 (often it's estimated as 3.14, but you want to make sure you have some extra decimal places to allow you to round off correctly at the end)

OpenStudy (anonymous):

did you actually get 63.617 (plus more digits)? That rounds to 63.62 cm squared

OpenStudy (anonymous):

that's what i got the first time but its says its not correct

OpenStudy (anonymous):

your class says "63.62 cm^2" is incorrect?

OpenStudy (anonymous):

you could try the answer listed earlier above... 63.58 but that's wrong :) He/She first rounded pi to 3.14, then solved to get 63.58. You should keep the digits in pi, then round the answer, not round the input values. But that doesn't mean your class solved it correctly... lots of math at this level wants you to estimate pi as 3.14. This problem doesn't tell you to do that though. So while 63.62 cm^2 is "most correct", it's possible your class answer is looking for 63.58 instead.

OpenStudy (anonymous):

then in that case i want to change my answer to 63.6428571428 or simply 63.64